D. Cramer’s Rule; 1. 6x+3y= 6; 5x-3y=3 2. 2x-3y=5; 4x+10y=4


1
Expert's answer
2022-03-16T14:18:20-0400

1.

"\begin{cases}
 6x+3y= 6 \\
 5x-3y=3
\end{cases}"


"\Delta = det\begin{vmatrix}
 6 & 3 \\
 5 & -3
\end{vmatrix} = 6\cdot(-3) - 3 \cdot 5 = -33"


"\Delta_1 = det\begin{vmatrix}
 6 & 3 \\
 3 & -3
\end{vmatrix} = 6 \cdot (-3) - 3 \cdot 3 = -27"


"\Delta_2 = det\begin{vmatrix}
 6 & 6 \\
 5 & 3
\end{vmatrix} = 6 \cdot 3 - 6 \cdot 5 = -12"


"x = \frac{\Delta_1}{\Delta} = \frac{-27}{-33} =\frac{9}{11}"


"y = \frac{\Delta_2}{\Delta} = \frac{-12}{-33} =\frac{4}{11}"


2.

"\begin{cases}
 2x-3y=5 \\
 4x+10y=4
\end{cases}"


"\Delta = det\begin{vmatrix}
 2 & -3 \\
 4 & 10
\end{vmatrix} = 2\cdot 10 - (-3) \cdot 4 = 32"


"\Delta_1 = det\begin{vmatrix}
 5 & -3 \\
 4 & 10
\end{vmatrix} = 5\cdot 10 - (-3) \cdot 4 = 62"


"\Delta_2 = det\begin{vmatrix}
 2 & 5 \\
 4 & 4
\end{vmatrix} = 2\cdot 4 - 5 \cdot 4 = -12"


"x = \frac{\Delta_1}{\Delta} = \frac{62}{32} = \frac{31}{16}"


"y = \frac{\Delta_2}{\Delta} = \frac{-12}{32} = \frac{-3}{8}"